OpenStudy (anonymous):

solve the quadratic equation by completing the square x^2+18x+76=0

OpenStudy (anonymous):

x^2 + 2.9.x +9^2 +76 -9^2=0 (x+9)^2=81-76 x+9=+-\[\sqrt{5}\] x=-9\[\pm \sqrt{5}\]